OpenStudy (anonymous):

How do you the rationalize the root in something like:

OpenStudy (anonymous):

\[\frac{ 5 }{ \sqrt{7} + 2}\]

OpenStudy (anonymous):

You multiply the numerator and the denominator by it's conjugate. The conjugate being \[\sqrt{7}-2\] So this is what you do: \[\frac{ 5 }{ \sqrt{7}+2 }(\frac{ \sqrt{7}-2 }{ \sqrt{7}-2 })\]

OpenStudy (anonymous):

Thank you (:

OpenStudy (anonymous):

You're welcome :)